OpenStudy (anonymous):

Solve and graph the absolute value inequality: |2x + 4| > 8. number line with open circles on −6 and 2, shading in between. number line with closed circles on −6 and 2, shading going in the opposite directions. number line with open circles on −6 and 2, shading going in the opposite directions. number line with open circles on −2 and 2, shading going in the opposite directions.

OpenStudy (anonymous):

@ganeshie8 @dumbcow @phi

OpenStudy (anonymous):

@Hero

OpenStudy (phi):

First, does the > mean open circle or closed circle ?

OpenStudy (anonymous):

open

OpenStudy (phi):

so cross off the choice that says "closed" that leaves 3 choices.

OpenStudy (anonymous):

I think it's A but I'm not sure

OpenStudy (phi):

how about the first choice number line with open circles on −6 and 2, shading in between. ? shading "in between" -6 and 2. Pick an easy number in between -6 and 2 and see if it "works" for example, let x=0 is |2 * 0 + 4 | > 8 ?

OpenStudy (phi):

what is 2 times 0 ? then add 4 what do you get

OpenStudy (anonymous):

the answer is false,

OpenStudy (anonymous):

2 * 0 = 0

OpenStudy (anonymous):

0 + 4 = 4

OpenStudy (anonymous):

4 is not bigger than 8

OpenStudy (phi):

and the | 4 | means make 4 positive. i.e. you get just 4 (because 4 is already positive) so cross off choice A.

OpenStudy (anonymous):

C? because D does not make sense to me..

OpenStudy (phi):

it's either C or D D says -4 should work (for example)

OpenStudy (phi):

if we test -4, we get | -4 * 2 + 4 | = | -8 + 4 | = | -4 | = 4 and 4 > 8 is false. so that leaves C

OpenStudy (anonymous):

Okay, Thanks! can you help me with one more?

OpenStudy (anonymous):

it is just like this one,

OpenStudy (phi):

the other way is to solve for x. we set up two relations 2x + 4 > 8 and -(2x + 4) > 8 solve the first one: add -4 to both sides 2x +4 - 4 > 8 -4 2x > 4 divide both sides by 2 x > 2 so x bigger than 2... shade in the numbers bigger than 2 now the 2nd relation -(2x + 4) > 8 distribute the - sign (it means multiply by -1) -2x - 4 > 8 add +4 to both sides -2x -4 + 4 > 8 +4 -2x > 12 now it is trick to multiply or divide by a negative number. To stay out of trouble, I ALWAYS ADD OR SUBTRACT add +2x to both sides. -2x + 2x > 2x+12 0 > 2x+12 now add -12 to both sides -12 > 2x divide both sides by 2 -6 > x or x< - 6 x is smaller than -6